OpenStudy (anonymous):

Solve on the interval 0 ≤ θ < 2π cos(2θ-π/2)=-1

terenzreignz (terenzreignz):

Don't I just love messing with trigonometric identities :D First of all... what is cos(-θ) equal to?

terenzreignz (terenzreignz):

Well, I'll let this one go... cos(-θ) = cos θ Agree?

OpenStudy (anonymous):

Yeah..

terenzreignz (terenzreignz):

So... \[\Large \cos\left(2\theta-\frac{\pi}{2}\right)=\cos\left[-\left(2\theta-\frac{\pi}{2}\right)\right]=\cos\left(\frac{\pi}{2}-2\theta\right)\]

OpenStudy (anonymous):

Wait why does it become negative?

terenzreignz (terenzreignz):

It didn't become negative.. I put in a negative sign. You did agree that cos(-x) = cos(x) So what I did was turn the 2θ-π/2 into π/2-2θ And because cos(-x) = cos(x) that doesn't change the value.

terenzreignz (terenzreignz):

The reason I did this is because we have another useful identity... \[\huge \cos\left(\frac{\pi}{2}-\alpha\right)=\sin(\alpha)\]

OpenStudy (anonymous):

Oh... haha okay

terenzreignz (terenzreignz):

So...\[\huge \huge \cos\left(2\theta-\frac{\pi}{2}\right)=\cos\left(\frac{\pi}{2}-2\theta\right)=\sin(2\theta)\] And I think you'll agree that that's much simpler.

OpenStudy (anonymous):

So it becomes sin(2θ)=-1?

terenzreignz (terenzreignz):

that it does :)